OpenStudy (anonymous):

Alright, guys, I'm confused. J + S = 150 2J + 3.5S = 405 Solving for J

hartnn (hartnn):

you have 2 methods : elimination, substitution. which one u wanna use ?

OpenStudy (anonymous):

I thought, S = 150 - J <-- sub 2J + 3.5(150 - J) = 405 2J + 525 - 3.5J = 405 -1.5J = -120 J = 80, is that right?

hartnn (hartnn):

yes, correct.

hartnn (hartnn):

any doubts

OpenStudy (anonymous):

No doubts.

hartnn (hartnn):

good :)

OpenStudy (anonymous):

@hartnn , got a question. If I had, 2x + y = 3, and I wanted to put in y = mx + b I would say: y = 2x + 3, or would the 2x be negative because I switched it's side?

hartnn (hartnn):

since u switched side its -2x

hartnn (hartnn):

a better way to explain that is 2x+y=3 subtracting 2x from both sides 2x-2x+y= -2x +3 i subtracted because we need to do inverse operation, since 2x was added to y, to isolate y, you subtract it.